OpenStudy (anonymous):

A man buys a motorcycle. After t months t value $V is given by V = 10000e^(-pt), where p is a constant. The value of the motorcycle after 12 months expected to be $4000. Calculate: 1) the expected value of the motorcycle after 18 months 2) the age of the motorcycle, to the nearest month, when its expected value will be $1000. WHAT I'VE DONE SO FAR: i) $10,000 ii) Do I find the value of P using the first given (i.e. price at 12 months) and just substitute t with 18 afterwards? iii) Do I find the value of P using the first given and just substitute V with $1000?

OpenStudy (anonymous):

ii) \[V = 10 000e ^{-pt}\] \[4000 = 10000e^{-12p}\] How do I apply natural log to this equation?

OpenStudy (anonymous):

@Aurora-Borealis, Thank you very much! :) I'm not actually sure how to apply this... I haven't seen this kind of method of answering before.

OpenStudy (anonymous):

Sorry my bad, it's wrong. :\

jimthompson5910 (jim_thompson5910):

V = 10000e^(-pt) 4000 = 10000e^(-p*12) 4000/10000 = e^(-12p) 0.4 = e^(-12p) ln(0.4) = -12p ln(0.4)/(-12) = p 0.07635756 = p p = 0.07635756 So the equation is V = 10000e^(-0.07635756t) ------------------------------------- a) Plug in t = 18 and evaluate to find its value after 18 months V = 10000e^(-0.07635756t) V = 10000e^(-0.07635756*18) V = 10000e^(-1.37443608) V = 10000(0.252982217) V = 2529.82217 So after 18 months, the value is $2,529.82

jimthompson5910 (jim_thompson5910):

b) Plug in V = 1000 and solve for t V = 10000e^(-0.07635756t) 1000 = 10000e^(-0.07635756t) 1000/10000 = e^(-0.07635756t) 0.1 = e^(-0.07635756t) ln(0.1) = -0.07635756t ln(0.1)/(-0.07635756) = t 30.1552995 = t t = 30.1552995 So it will take roughly 30.1552995 months for the value to become $1,000

OpenStudy (anonymous):

Thank you very, very much! :D

jimthompson5910 (jim_thompson5910):

you're welcome